OpenStudy (anonymous):

Find the equations of the tangent and normal to the following curves at the point indicated. x²-2xy-2y² = 1; (-3, -1) Help please :)

OpenStudy (anonymous):

i think its like this.... not so sure uh... since y' = slope = m find y' : 2x-2(xy'+y)=0 2x-2xy'-2y=0 x-xy'-y=0 y'=(x-y)/x for tangent : y+1=(x-y)/x * (x+3) for normal: y+1=-x/(x-y) * (x+3)

OpenStudy (anonymous):

Alright, thanks. I'm gonna figure it too. :)

OpenStudy (nikvist):

\[x^2-2xy-2y^2=1\]\[(x^2-2xy-2y^2)'=0\]\[2x-2y-2xy'-4yy'=0\]\[x-y=(x+2y)y'\]\[y'=\frac{x-y}{x+2y}\]\[y'(-3,-1)=\frac{-3-(-1)}{-3+2(-1)}=\frac{-2}{-5}=\frac{2}{5}\] tangent: \[y+1=\frac{2}{5}(x+3)\] normal: \[y+1=-\frac{5}{2}(x+3)\]